### v1

# Traffic Scene Generation from Natural Language Description for Autonomous Vehicles with Large Language Model

Abstract
Text-to-scene generation, transforming textual descriptions into detailed scenes, typically relies on generating key scenarios along predetermined paths, constraining environmental diversity and limiting customization flexibility. To address these limitations, we propose a novel text-to-traffic scene framework that leverages a large language model to generate diverse traffic scenarios within the Carla simulator based on natural language descriptions. Users can define specific parameters such as weather conditions, vehicle types, and road signals, while our pipeline can autonomously select the starting point and scenario details, generating scenes from scratch without relying on predetermined locations or trajectories. Furthermore, our framework supports both critical and routine traffic scenarios, enhancing its applicability. Experimental results indicate that our approach promotes diverse agent planning and road selection, enhancing the training of autonomous agents in traffic environments. Notably, our methodology has achieved a 16% reduction in average collision rates. Our work is made publicly available at https://basiclab.github.io/TTSG/.
I Introduction
In autonomous systems, the ability of agents to effectively interact with their surroundings, recognize hazardous objects in the environment and perform safe maneuvers enhances the accuracy of traffic trajectory prediction and planning. To achieve these tasks, datasets such as nuScenes [1, 2] and Waymo [3, 4] are made to provide substantial real-world data, aiding researchers to train a model that can handle driving conditions under several environments and implicitly recognizing different driving rules. Nevertheless, the existing datasets often lack variability in critical conditions or scene types, and it can be hard for agents to interact with the environment since conducting diverse real-world tests with autonomous agents is costly and poses safety risks. Consequently, there is increasing reliance on simulated environments [5, 6] to gather data and evaluate agents. These simulations, however, typically require predefined routes and spawning positions for agents, which limits testing flexibility.
We aim to generate traffic scenes through natural language inputs, enabling dynamic and flexible scenario creation. For instance, given an input prompt: “Schedule a daily traffic at the intersection”, we can obtain a traffic scenario that matches the description. Recent advancements have highlighted the powerful capabilities of Large Language Models (LLMs) in reasoning [7, 8], planning [9, 10], and even code generation [11]. One existing approach designed specifically for generating critical scenarios, ChatScene [12], leverages an LLM to generate critical conditions within the Carla environment [5]. It does so by first selecting a few predefined positions and agents to perform the scenario and adopting an LLM agent to perform in-context learning with a few code examples to generate code for controlling agents. However, while ChatScene uses predefined positions to reduce the complexity of generating scenarios, it sacrifices diversity. Furthermore, the reliance on few-shot examples and fixed formats restrict customization options, such as agent types, road signals, and objects, limiting the creation of tailored traffic scenarios.
Figure 1: Comparison between ChatScene (top) and ours (bottom).
To increase diversity and allow more customized options, we introduce a novel pipeline that can directly provide the agent scheduling, including types, actions, positions, and road conditions, and can automatically select proper positions for scene generation without requiring human-specified trajectories or spawning positions. The comparison of ChatScene and our pipeline is shown in Fig. 1. Specifically, our system first uses the LLM agent to perform the prompt analysis to decompose different road conditions and required agents. The potential roads that match the condition are retrieved from the database. Then, our system prompts the LLM agent to provide the agent planning that should be created within the scene based on the analysis results. Since agents can take different actions, we should choose the road that allows them to perform specific actions. For instance, there should be a left road if an agent is going to turn left. Therefore, we rank the potential roads retrieved before with the agent’s plans to find the highest one for scene generation. Finally, we design our rendering interface to translate the planning results from LLM into the traffic scenario. Since our approach can automatically select the road for spawning and place the agent based on their plans, no predefined selection of points and trajectories is required. We also build an agent database that includes different types of agents and actions that enable users to choose from.
We demonstrate the efficacy of our framework by training an agent on our generated critical scenarios and evaluating the generation quality under different component designs. Our contributions are summarized as follows:
  • We propose a novel system for generating traffic scenes from natural language descriptions through road retrieval and agent planning pipeline with an LLM.
  • Our system supports the diverse generation of traffic scenes on road selection and agent planning and allows customized options for different scenario usages.
  • Our approach can be used to train an agent under critical scenarios and obtain a 16% reduction in collision rate.
II Related Work
II-A Traffic Scene Generation
Traffic scene generation using generative models has captured significant attention in recent research. Approaches such as [13, 14, 15] have leveraged learning-based methods to schedule agent trajectories within given scenario maps. To imitate similar traffic scenes in a new place, RealGen [16] employs Retrieval Augmented Generation (RAG) [17] technique to enhance scene generation by retrieving existing scenarios as a prior for aiding the decoder in producing relevant traffic scenarios. As different platforms and datasets have sprung up, ScenarioNet [18] introduces a unified traffic simulation platform that combines synthetic and real-world scenarios to enhance large-scale multi-agent learning and explore adversarial attacks.
A recent line of work focuses on critical scenario generation as it poses significant risks during driving [19, 20, 21]. These techniques often adopt adversarial training under scenarios initially generated in critical states to ensure that driving agents can manage critical cases and enhance safety. To generate the critical scenario automatically with text, ChatScene [12] facilitates text-to-critical-scenario conversion using an LLM agent that employs the Scenic programming language [22] to manage autonomous agents. Unlike previous work, we implement a road-and-agent retrieval system that facilitates text-to-scene conversions applicable to diverse scenarios, including multi-agent driving and critical scenario production. This approach allows for greater flexibility compared to ChatScene, which necessitates predefined spawning points and offers limited customized options by enabling users to generate scenes directly from text without the need to select maps or locations.
II-B Autonomous Driving with LLM
Large Language Models (LLMs), such as GPT series [23] and Llama [24], have demonstrated substantial capabilities in natural language processing and have been increasingly adapted for diverse applications [25, 26, 27]. In the domain of autonomous driving, one key objective is to enable decision-making based on robust reasoning. Due to their potent planning and reasoning capabilities, LLMs have been integrated into vision language model (VLM) systems for trajectory scheduling and identifying safety-critical objects.
For instance, ADAPT [7] utilizes LLMs for video captioning and reasoning of the driving decisions. Regarding action planning, DriveLM [9] and DriveLLM [28] employ LLMs to dynamically plan vehicle actions based on image inputs. To incorporate different modalities such as point clouds, DriveMLM [10] apply Q-former [29] to translate all inputs into tokens that can be read by an LLM. To deploy the agent into a continuous driving environment, DiLu [30] incorporates a memory module within its framework, enabling the continuous learning and application of new knowledge derived from environmental interactions. Additionally, RAG-Driver [31] implements a RAG system to analogize current driving conditions with historical scenarios, thereby providing precise planning inputs with past experience.
As these works show that LLM agents are suitable for reasoning and planning for driving scenarios, we aim to combine the LLM to generate the traffic scene from text inputs. This combination of LLM agents allows us to translate text input into proper scene scheduling by analyzing the input scenario to decide road conditions and provide proper agent planning.
III Traffic Scene Generation
Figure 2: Illustration of the TTSG pipeline, detailing five principal stages: analysis, road candidate retrieval, agent planning, road ranking, and generation. The analysis stage decompose the user’s input to refine the processing in subsequent phases. This is followed by an LLM agent determining retrieval conditions to identify appropriate road candidates and facilitating precise planning for agents, which includes types, positions, and actions. In cases where multiple road options are retrieved, a ranking strategy prioritizes roads that best match the desired conditions and offer sufficient lanes and space for scenario depiction. The traffic scene is ultimately generated through our custom rendering interface, utilizing the selected road and agent details.
In this section, we detail the development of our Text-to-Traffic Scene Generation (TTSG) framework, which leverages an LLM for planning and retrieval to generate traffic scenes from natural language descriptions. Our methodology starts with constructing a database that categorizes the attributes of roads and agents. Given a scene description as shown in Fig. 2-(a), an LLM agent first analyzes the input prompts, determining the road conditions for retrieval and devising a plan for agent scheduling. Following this, the potential road candidates undergo a ranking process, which evaluates the properties and actions of agents to identify the most appropriate selection for scene generation. This ranking strategy is crucial for ensuring the optimal match of road properties to the intended scenario. The characteristics of the selected road are then integrated with the agents’ planning outcomes and visualized using our custom rendering interface. To validate our pipeline’s effectiveness and conduct thorough evaluations, the Carla simulator is utilized as our primary demonstration platform. For effective record--keeping and system integration, the outputs from the LLM agent are all stored in JSON format.
III-A Road and Agent Database Construction
Constructing a road database is crucial for augmenting the capabilities of the large language model (LLM) agent. This database is designed to provide spawn information, enabling the LLM to deliver precise planning results even for scenarios it has not previously encountered. Unlike previous work [12], which focuses on gathering code snippets, our database categorizes roads, including traffic signals and road objects found within each road segment. This strategy allows for dynamic scenario generation without the need for predefined geometries and spawning positions, enhancing the flexibility of our system across various traffic conditions.
Specifically, to construct this road database for LLM, we convert each map into the OpenDRIVE format—an open-source XML syntax that standardizes road network descriptions. [^1] https://www.asam.net/standards/detail/opendrive/ We parse details such as signals, objects, junctions, and lane information and organize them into a graph structure. In this structure, edges represent road connectivity, which simplifies the querying process. Additionally, agents are categorized to enable targeted queries based on agent type, enhancing the system’s usability and effectiveness.
III-B Text-to-Traffic Scene
After establishing the road and agent database, we adopt an LLM agent and design the system prompt to utilize user inputs for retrieving appropriate roads and agents, thereby generating the desired traffic scenarios with accurate planning. The process, depicted in Fig. 2, consists of analysis, retrieval, planning, ranking, and generation. To ensure seamless operation, we conduct format verification on the output to confirm that the keys, types, and values after each step are compatible with our pipeline. Should any discrepancies arise during this validation process, the original input and a detailed error message are resubmitted to the LLM agent for correction, thereby guaranteeing the accuracy and feasibility of the generated traffic scenarios.
III-B1 Analysis
The primary objective in the analysis mode is to equip the LLM agent with a complete understanding of the user’s requirements, facilitating precise planning decisions. One promising approach is to leverage the chain-of-thought method [32, 33], which involves generating step-by-step explanations. However, this approach requires more output tokens from the LLM, which can be computationally expensive. To streamline the process, we instead aim to provide the LLM agent with enriched contextual knowledge by first decomposing each component, including the required signals, objects, and agents from the user’s input as illustrated in Fig. 2-(a). This decomposed output aids the LLM in processing and interpreting the user’s natural language description more effectively, setting the stage for a precise understanding of the input for scenario generation.
III-B2 Road Candidate Retrieval
This mode focuses on querying the road database to find the road that aligns with the user’s specifications and ensures accurate scene generation. The retrieval process is crucial for selecting optimal road locations and obtaining essential information for scenario planning. Leveraging the output from the analysis phase and the user’s prompt, our system efficiently narrows the search to identify roads within the road database that meet specific criteria, including the number of lanes, having or not having any specific traffic signals, and the roadside objects (e.g., speed signs or stop signs).
III-B3 Agent Planning
The planning mode is designed to provide the correct agent setup based on the user inputs. For instance, given the input “two cars in front block the ego car” and the analysis results of the same prompt to the LLM agent, it should provide a plan that ensures: 1) two agents, identified as cars, are positioned in front of the ego car, 2) within the same lane, and 3) their actions are programmed to either stop or remain idle. This strategic placement and behavior setup facilitate realistic and targeted scenario simulations. Additionally, each agent can be oriented in one of eight directions relative to the ego agent or even positioned on adjacent roads, enhancing the dynamic and realism of the scenario. Complex behaviors like aggressive driving styles are configurable by adding the adjective to the agent such as “a dangerous cyclist”.
Figure 3: Demonstration of multi-agent planning with prompts detailed in each sub-caption.
Figure 4: Demonstration of critical scenarios with prompts detailed in each sub-caption.
III-B4 Road Ranking
Upon identifying potential road candidates at the road candidate retrieval step, our pipeline faces a decision point: selecting a road randomly or employing a more systematic approach. We opt for the latter since we can implement a ranking strategy, which refines the selection process to ensure the most suitable road is chosen for the scenario. This involves scoring each road based on the agent plan, such as the directional options available to the agent (e.g., turning left, right, or continuing straight), the suitability of the road type for agent deployment (e.g., driving lanes, shoulder, or sidewalk), and the maximum capacity of a road. The ranking process enhances our confidence that the highest-scoring roads properly match the input’s requirements since it helps consider the agent’s planning, making the selected road suitable for an agent to perform different actions. Finally, we adopt a random selection to pick up a road from multiple roads with the same highest score. This ranking-first-with-random-later strategy ensures accurate and diverse road picking for a generation.
III-B5 Scene Generation
Finally, our pipeline integrates road attributes and agent plans within our rendering interface to generate traffic scenes. Our rendering interface is a function that can translate the JSON outputs from an LLM agent into the traffic scene within Carla. It first locates the selected road, spawns the agents, and sets their properties according to the planning. The interface will also control the agent’s action during the generation process to ensure the output scenes match the planning settings. The recorded images of the ego agent and the trajectories of all participating agents can be applied for subsequent training or analysis.
III-C Applications under Different Scenarios
III-C1 Multi-Agent Planning
Our pipeline can be useful in training models that learn to determine trajectories or actions for controlling one or multiple agents simultaneously. Furthermore, users can customize these scenarios by setting them under varying weather conditions, such as during a cloudy day or a rainy night. Another application is simulating the condition with emergency vehicles, enabling scenarios where ambulances, fire trucks, or police vehicles navigate without impediments, thereby enhancing the realism and utility of the training data. We provide examples in Fig. 3.
III-C2 Critical Scenario
A significant aspect of our scene generation capability focuses on critical scenarios, which are essential for testing the robustness of autonomous driving systems. Our system can be prompted to create scenarios with potentially dangerous settings. For example, in Fig. 4(a), we can generate scenes where visibility is obstructed given input presented in the sub-caption, causing pedestrians and vehicles to be unseen by each other. Additionally, with the road conditions, we can create an unprotected left-turn scene with no traffic light and stop sign as presented in Fig. 4(b).
IV Experiments
IV-A Pipeline Setup
In the experimental setup, we utilize the Carla simulator, version 0.9.15, and employ the built-in maps from Town01 to Town07 to construct our road database. Roads lacking sufficient space for agent deployment are excluded. Our agent database accommodates nine distinct types of agents: ambulance, police car, firetruck, bus, truck, motorcycle, cyclist, car, and pedestrian. The simulation engine supports six actions: turning, lane changing, straight, stop, crossing road, and blocking, with denoting optional left or right directions. Additionally, as for position relative to the ego agent, our system supports twelve directional positions, including eight basic compass directions and specific placements like left road, right road, opposite straight road, and near the ego destination. We provide the illustration for the action and the relative position in Fig. 5. As for the LLM agent, we adopt the GPT-4o [23] to achieve analysis, road candidate retrieval, and agent planning.
IV-B Generation for Critical Scenarios
To assess the efficacy of our framework to train an autonomous agent, we examine our approach to critical scenarios and utilize SafeBench [34], a benchmark designed for the safety evaluation of autonomous vehicle systems. Our pipeline is employed to create scenes across three distinct challenging safety scenarios selected from ChatScene [12], aiming to conduct adversarial training to bolster the ego vehicle’s safe-driving skills. We prompt the LLM agent with the scenario descriptions and use the planning results from the LLM and our rendering interface to train the ego agents under a controlled policy using a soft-actor-critic model [35]. During training, the ego agent is controlled by a learnable policy model, and other agents that create the scenario are controlled by our engine.
For consistency and comparability with the studies reported in [12], the vehicles are evaluated in the same environment of [12] to measure their Collision Rate (CR) and Overall Score (OS), which includes various metrics of driving performance. The outcomes are presented in Table I, and each column is colored as best and second best.
In this table, we compare our results with Learning-to-Collide (LS) [36], AdvSim (AS) [37], Adversarial Trajectory Optimization (AT) [38], and ChatScene (CS) [12]. Our findings indicate that our text-to-scene pipeline effectively trains agents to handle critical conditions, thereby enhancing safety and reducing collision risks. Notably, our agents are proficient at managing obstacles and yielding to oncoming cars, thus avoiding dangerous collisions in the first and third scenario types. Despite achieving a higher collision rate and a lower driving score in the second scenario, where agents occasionally fail to decelerate when another vehicle attempts a risky maneuver, our system demonstrates competence in most cases to deal with the dangerous lane-changing scenarios, surpassing the majority of comparative approaches. Overall, our approach successfully prepares agents for critical scenarios, enabling them to learn from these experiences and elevate their safety capabilities.
Figure 5: Illustration of relative position (left) and the action (right).
TABLE I: Results evaluated on Critical Scenario.
Algo. Metric Scenario Avg.
Straight Obstacle Lane Changing Unprotected Left-turn
[…]
AS 0.230 0.530 0.050 0.270
AT 0.140 0.300 0.000 0.150
CS 0.030 0.110 0.100 0.080
Ours 0.000 0.200 0.000 0.067
[…]
AS 0.784 0.666 0.937 0.796
AT 0.849 0.803 0.948 0.867
CS 0.905 0.906 0.903 0.905
ours 0.955 0.861 0.951 0.922
TABLE II: Evaluation Results of Scene Diversity.
Metric Scenario Avg.
Normal Critical Conditional
Daily Traffic Intersection Pedestrian Crushing Blocking Agent Dangerous Cut-off Only having Two-wheel Vehicles Having Emergency Vehicles Rainy Weather
[…]
IV-C Ablation Study
The primary objective of our framework design is to provide a reliable pipeline to generate traffic scenes. Therefore, we evaluate the efficacy of different components by adopting ten distinct prompts under diverse scenarios. The scenarios encompass three normal, four critical, and three specific road conditions, such as the presence of traffic lights or the absence of crossroads or a stop sign on the road.
TABLE III: Analysis of pipeline design.
Quality Scenario Avg.
Normal Critical Conditional
[…]
w/o. analysis .917 .667 .833 .750 .750 .917 .833 .775
w./o analysis .917 1.00 .875 .750 1.00 .917 .925 .875
Scene Quality $\uparrow TM\uparrow$ $\uparrow TM\uparrow$
w/o. ranking 0.667 0.450 0.600 10.560
w./o ranking 0.867 0.750 0.800 0.800
IV-C1 Plan Quality
We assess plan quality by measuring the accuracy of planning results from the LLM agent using two metrics: Agent Accuracy (AA) and Road Accuracy (RA). AA evaluates the correctness of agent attributes, including position, action, and type. RA assesses the accuracy of the road conditions, such as signals, objects, and lane counts. As indicated in the first part of Table III, incorporating an analysis step into the reasoning process significantly enhances performance. This is because the analysis step provides the agent with more contextual information, leading to more precise decisions. Crucially, integrating an analysis step helps stabilize the number of agents, significantly mitigating the issue of generating unnecessary agents that substantially detract from the agent’s accuracy.
IV-C2 Scene Quality
To evaluate the impact of our road ranking strategy on scene quality, we use Text Matching (TM) to examine the correctness between the prompt and the generated scene. To ensure fairness and isolate the effect of the ranking strategy, identical road conditions and agent planning are maintained across evaluations, and each scenario is tested five times to calculate average outcomes. The results are detailed in the second part of Table III that the ranking approach consistently shows superior performance. The improvement is largely because, without ranking, the selection of roads is random, often disregarding crucial factors such as turn permissions or the adequacy of spawning points. Conversely, the road ranking strategy ensures the selection of roads that optimally align with the specified agent planning, thus enhancing the fidelity of the generated traffic scenes to the input prompts.
IV-D Diversity Test
A key claim of our approach is its ability to generate varied traffic scenes without reliance on fixed selections or predefined spawning points. We construct eight distinct scenarios under three main categories, normal, critical, and conditional, to validate the diversity of our generated results. To examine our pipeline’s ability to generate proper and diverse results, we provide no clues about the roads and agents within the text input. Each scenario is initiated with a text prompt formatted as: “Please create a scene for <scenario>” to generate traffic scenes.
We assess the diversity of these scenarios by generating each scene five times and compute the metrics, including Agent Diversity (AD) and Road Diversity (RD). These diversity metrics are calculated as the ratio of unique objects to the total number of objects. For agents, diversity considers variations in agent type, action, and relative position; for roads, diversity is assessed by unique road IDs (different directions on the same road is considered the same). Additionally, to verify the practicality and accuracy of our rendering interface, we include the metric Text Matching (TM), which evaluates whether the generated scenes accurately reflect the input prompts using a binary matched/unmatched criterion. The comprehensive results are detailed in Table II.
As the table indicates, our pipeline consistently provides a nearly unique road selection for each scenario. For the AD score, we observe lower scores in scenarios with specific types such as pedestrian crushing and emergency vehicles. This outcome is anticipated as more detailed agent descriptions naturally lead to similar planning results. Conversely, scenarios with less specific cues, such as daily traffic or intersection under normal scenarios, blocking agent under critical scenarios, or rainy weather under conditional scenarios, achieve higher diversity scores. These results demonstrate that our approach effectively leverages the LLM agent to generate diverse road and agent planning. Regarding the TM scores, despite lower scores in scenarios that involve precise timing (pedestrian crushing) or have restrictive conditions (only having two-wheel vehicles), our rendering interface accurately produce the majority of the scenarios, allowing us to obtain precise traffic scenes from text inputs.
V Conclusion and Future Work
In this paper, we introduce a novel framework that leverages natural language descriptions for traffic scene generation facilitated by a Large Language Model. Our pipeline is able to automatically identify diverse and optimal locations for scene generation and allows customized options. Our pipeline starts by analyzing user input to decompose the requirements, performing road candidate retrieval, and planning detailed agents. To ensure the selected scene is suitable for an agent to act and be spawned, we adopt a road ranking strategy to match all the conditions. We provide experiments on diversity tests and adversarial training on the critical scenes to examine our pipeline ability to generate diverse and reliable scenes. Looking ahead, we plan to enhance the system’s practicality for agent and road generation to enable the creation of non-existent signals and objects and design a generated model that can transfer simulated scenes into a real-world video. This will further expand our framework’s ability to create a real and dynamic scene.