### v2

# Traffic Scene Generation from Natural Language Description for Autonomous Vehicles with Large Language Model

Abstract
Text-to-scene generation typically limits environmental diversity by generating key scenarios along predetermined paths. To address these constraints, we propose a novel text-to-traffic scene framework that leverages a large language model (LLM) to autonomously generate diverse traffic scenarios for the CARLA simulator based on natural language descriptions. Our pipeline comprises several key stages: (1) Prompt Analysis, where natural language inputs are decomposed; (2) Road Retrieval, selecting optimal roads from a database; (3) Agent Planning, detailing agent types and behaviors; (4) Road Ranking, scoring roads to match scenario requirements; and (5) Scene Generation, rendering the planned scenarios in the simulator. This framework supports both routine and critical traffic scenarios, enhancing its applicability. We demonstrate that our approach not only diversifies agent planning and road selection but also significantly reduces the average collision rate from 8% to 3.5% in SafeBench. Additionally, our framework improves narration and reasoning for driving captioning tasks. Our contributions and resources are publicly available at https://basiclab.github.io/TTSG/.
I Introduction
In autonomous systems, motion planning models often require a lot of data training to earn robust driving capabilities. An interactive training environment that can generate scenarios based on instructions can greatly accelerate model training and reduce reliance on large amounts of data. Previous work on datasets, such as nuScenes [1, 2] and Waymo [3, 4] are made to provide substantial real-world data, aiding researchers to train a model that can handle driving conditions under several environments and implicitly recognizing different driving rules. Nevertheless, the existing datasets often lack variability in critical conditions or scene types, and it can be hard for agents to interact with the environment since conducting diverse real-world tests with autonomous agents is costly and poses safety risks. Consequently, there is increasing reliance on simulated environments [5, 6] to gather data and evaluate agents. These simulations, however, typically require predefined routes and spawning positions for agents, which limits testing flexibility.
We aim to generate traffic scenes through natural language inputs, enabling dynamic and flexible scenario creation. For instance, given an input prompt: “Schedule a daily traffic at the intersection”, we can obtain a traffic scenario that matches the description. Recent advancements have highlighted the powerful capabilities of Large Language Models (LLMs) in reasoning [7, 8], planning [9, 10], and even code generation [11]. A previous approach, ChatScene [12], has proposed to generate specific critical scenarios in CARLA [5]. However, as [12] uses predefined positions to reduce the complexity of generating scenarios, it sacrifices diversity. Furthermore, the reliance on few-shot examples and fixed formats restrict customization options, such as agent types, road signals, and objects, limiting the creation of tailored traffic scenarios.
Figure 1: Simplified pipeline of our approach. Our approach takes the scene description from the users and provides the correspondence traffic scene.
In this work, we introduce a novel pipeline that can directly provide the agent scheduling, including types, actions, positions, and road conditions, and can automatically select proper positions for scene generation without requiring human-specified spawning positions. The simplified pipeline of our framework is shown in Fig. 1. Specifically, our framework first uses the LLM agent to perform the prompt analysis to decompose different road conditions and required agents. The potential roads that match the condition are retrieved from the database. Then, our framework prompts the LLM agent to provide the agent planning that should be created within the scene based on the analysis results. Since agents can take different actions, the pipeline should choose the road that allows the agents to perform specific actions. For instance, there should be a left road if an agent is going to turn left. Therefore, we insert a ranking mechanism in the pipeline to score potential roads by counting the achieved requirements of the agent’s plans to find the highest one for scene generation. Finally, we design our rendering interface to translate the planning results from LLM into the traffic scenario. Since our approach can automatically select the road for spawning and place the agent based on their plans, no predefined selection of points and maps is required. We also build an agent database that includes different types of agents and actions that enable users to choose from.
We demonstrate the efficacy of our framework by training an agent on our generated critical scenarios for both navigating and captioning tasks. To study the importance of each component, we evaluate the generation quality under different designs to provide a deeper understanding of our pipeline. Our contributions are summarized as follows:
  • We propose a novel framework for generating traffic scenes from natural language descriptions through road retrieval and agent planning pipeline with an LLM.
  • Our framework supports the diverse generation of traffic scenes on road selection and agent planning and allows customized options for different scenario usages.
  • Our approach can be used to train an agent under critical scenarios to reduce collision rate and to improve driving score and can also be incorporated to generate data for training a captioning model to narrate and reason the actions for critical scenarios.
II Related Work
II-A Traffic Scene Generation
Traffic scene generation using generative models has captured significant attention in recent research. Approaches such as [13, 14, 15] have leveraged learning-based methods to schedule agent trajectories within given scenario maps. To imitate similar traffic scenes in a new place, RealGen [16] employs Retrieval Augmented Generation (RAG) [17] technique to enhance scene generation by retrieving existing scenarios as a prior for aiding the decoder in producing relevant traffic scenarios. As different platforms and datasets have sprung up, ScenarioNet [18] introduces a unified traffic simulation platform that combines synthetic and real-world scenarios to enhance large-scale multi-agent learning and explore adversarial attacks.
A recent line of work focuses on critical scenario generation as it poses significant risks during driving [19, 20, 21]. These techniques often adopt adversarial training under scenarios initially generated in critical states to ensure that driving agents can manage critical cases and enhance safety. To generate the critical scenario automatically with text, ChatScene [12] facilitates text-to-critical-scenario conversion using an LLM agent that employs the Scenic programming language [22] to manage autonomous agents. Unlike previous work, we implement a road-and-agent retrieval method that facilitates text-to-scene conversions applicable to diverse scenarios, including multi-agent driving and critical scenario production. This approach allows for greater flexibility without necessitating predefined spawning points and offers various customized options by enabling users to generate scenes directly from text.
II-B Autonomous Driving with LLM
Large Language Models (LLMs), such as GPT series [23] and Llama [24], have demonstrated substantial capabilities in natural language processing and have been increasingly adapted for diverse applications [25, 26, 27, 28]. In the domain of autonomous driving, one key objective is to enable decision-making based on robust reasoning. Due to their potent planning and reasoning capabilities, LLMs have been integrated into vision language models (VLMs) for trajectory scheduling and identifying safety-critical objects.
For instance, ADAPT [7], and DriveGPT4 [29] utilize LLMs for video captioning and reasoning of the driving decisions. Regarding action planning, DriveLM [9] and DriveLLM [30] employ LLMs to dynamically plan vehicle actions based on image inputs. To incorporate different modalities such as point clouds, DriveMLM [10] apply Q-former [31] to translate all inputs into tokens that can be read by an LLM. To deploy the agent into a continuous driving environment, DiLu [32] incorporates a memory module within its framework, enabling the continuous learning and application of new knowledge derived from environmental interactions. Additionally, RAG-Driver [33] implements a RAG system to analogize current driving conditions with historical scenarios, thereby providing precise planning inputs with past experience.
As these works show that LLM agents are suitable for reasoning and planning for driving scenarios, we aim to combine the LLM to generate the traffic scene from text inputs. This combination of LLM agents allows us to translate text input into proper scene scheduling by analyzing the input scenario to decide road conditions and provide proper agent planning.
III Traffic Scene Generation
Figure 2: Illustration of the TTSG pipeline, detailing five principal stages: analysis, road candidate retrieval, agent planning, road ranking, and generation. The analysis stage decomposes the user’s input to refine the processing in subsequent phases. This is followed by an LLM agent determining retrieval conditions to identify appropriate road candidates and facilitating precise planning for agents, which includes types, positions, and actions. In cases where multiple road options are retrieved, a ranking strategy prioritizes roads that best match the desired conditions and offer sufficient lanes and space for scenario depiction. The traffic scene is ultimately generated through our custom rendering interface, utilizing the selected road and agent details.
In this section, we detail the development of our Text-to-Traffic Scene Generation (TTSG) framework, which leverages an LLM for planning and retrieval to generate traffic scenes from natural language descriptions. Our methodology starts with constructing a database that categorizes the attributes of roads and agents (Sec. III-A). Given a scene description as shown in Fig. 2(a), an LLM agent first analyzes the input prompts, determining the road conditions for retrieval and devising a plan for agent scheduling (Sec. III-B1, III-B2 and III-B3). Following this, the potential road candidates undergo a ranking process, which evaluates the properties and actions of agents to identify the most appropriate selection for scene generation (Sec. III-B4). This ranking strategy is crucial for ensuring the optimal match of road properties to the intended scenario. The characteristics of the selected road are then integrated with the agents’ planning outcomes and visualized using our custom rendering interface. To validate our pipeline’s effectiveness and conduct thorough evaluations, the CARLA simulator is utilized as our primary demonstration platform. For effective record--keeping and system integration, the outputs from the LLM agent are all stored in JSON format.
III-A Road and Agent Database Construction
Constructing a road database is crucial for augmenting the capabilities of the large language model (LLM) agent. This database is designed to provide spawn information, enabling the LLM to deliver precise planning results even for scenarios it has not previously encountered. Unlike previous work [12], which focuses on gathering code snippets, our database categorizes roads, including traffic signals and road objects found within each road segment. This strategy allows for dynamic scenario generation without the need for predefined geometries and spawning positions, enhancing the flexibility of our framework across various traffic conditions.
Specifically, to construct this road database for LLM, we convert each map into the OpenDRIVE format—an open-source XML syntax that standardizes road network descriptions. [^1] https://www.asam.net/standards/detail/opendrive/ We parse details such as signals, objects, junctions, and lane information and organize them into a graph structure. In this structure, edges represent road connectivity, which simplifies the querying process. Additionally, agents are categorized to enable targeted queries based on agent type, enhancing the system’s usability and effectiveness.
III-B Text-to-Traffic Scene
After establishing the road and agent database, we adopt an LLM agent and design the system prompt to utilize user inputs for retrieving appropriate roads and agents, thereby generating the desired traffic scenarios with accurate planning. The process, depicted in Fig. 2, consists of analysis, retrieval, planning, ranking, and generation. To ensure seamless operation, we conduct format verification on the output to confirm that the keys, types, and values after each step are compatible with our pipeline. Should any discrepancies arise during this validation process, the original input and a detailed error message are resubmitted to the LLM agent for correction, thereby guaranteeing the accuracy and feasibility of the generated traffic scenarios.
III-B1 Analysis
The primary objective in the analysis mode is to equip the LLM agent with a complete understanding of the user’s requirements, facilitating precise planning decisions. One promising approach is to leverage the chain-of-thought method [34, 35], which involves generating step-by-step explanations. However, this approach requires more output tokens from the LLM, which can be computationally expensive. To streamline the process, we instead aim to provide the LLM agent with enriched contextual knowledge by first decomposing each component, including the required signals, objects, and agents from the user’s input as illustrated in Fig. 2(a). This decomposed output aids the LLM in processing and interpreting the user’s natural language description more effectively, setting the stage for a precise understanding of the input for scenario generation.
III-B2 Road Candidate Retrieval
This mode focuses on querying the road database to find the road that aligns with the user’s specifications and ensures accurate scene generation. The retrieval process is crucial for selecting optimal road locations and obtaining essential information for scenario planning. Leveraging the output from the analysis phase and the user’s prompt, our framework efficiently narrows the search to identify roads within the road database that meet specific criteria, including the number of lanes, having or not having any specific traffic signals, and the roadside objects (e.g., speed signs or stop signs).
III-B3 Agent Planning
The planning mode is designed to provide the correct agent setup based on the user inputs. For instance, given the input “two cars in front block the ego car” and the analysis results of the same prompt to the LLM agent, it should provide a plan that ensures: 1) two agents, identified as cars, are positioned in front of the ego car, 2) within the same lane, and 3) their actions are programmed to either stop or remain idle. This strategic placement and behavior setup facilitate realistic and targeted scenario simulations. Additionally, each agent can be oriented in one of eight directions relative to the ego agent or even positioned on adjacent roads, enhancing the dynamic and realism of the scenario. Complex behaviors like aggressive driving styles are configurable by adding the adjective to the agent such as “a dangerous cyclist”.
Figure 3: Demonstration of multi-agent planning with prompts detailed in each sub-caption.
Figure 4: Demonstration of critical scenarios with prompts detailed in each sub-caption.
III-B4 Road Ranking
Upon identifying potential road candidates at the road candidate retrieval step, our pipeline faces a decision point: selecting a road randomly or employing a more systematic approach. We opt for the latter since we can implement a ranking strategy, which refines the selection process to ensure the most suitable road is chosen for the scenario. This involves scoring each road based on the agent plan, such as the directional options available to the agent (e.g., turning left, right, or continuing straight), the suitability of the road type for agent deployment (e.g., driving lanes, shoulder, or sidewalk), and the maximum capacity of a road. The ranking process enhances our confidence that the highest-scoring roads properly match the input’s requirements since it helps consider the agent’s planning, making the selected road suitable for an agent to perform different actions. Finally, we adopt a random selection to pick up a road from multiple roads with the same highest score. This ranking-first-with-random-later strategy ensures accurate and diverse road picking for a generation.
III-B5 Scene Generation
Finally, our pipeline integrates road attributes and agent plans within our rendering interface to generate traffic scenes. Our rendering interface is a function that can translate the JSON outputs from an LLM agent into the traffic scene within CARLA. It first locates the selected road, spawns the agents, and sets their properties according to the planning. The interface will also control the agent’s action during the generation process to ensure the output scenes match the planning settings. The recorded images of the ego agent and the trajectories of all participating agents can be applied for subsequent training or analysis.
TABLE I: Results evaluated on Critical Scenario with SafeBench.
[…]
Straight Obstacle Lane Changing Unprotected Left-turn Average Straight Obstacle Lane Changing Unprotected Left-turn Average
LC [36] IROS’20 0.120 0.510 0.000 0.210 0.827 0.684 0.954 0.822
AS [37] CVPR’21 0.230 0.430 0.050 0.270 0.784 0.666 0.937 0.796
AT [38] CVPR’22 0.140 0.300 0.000 0.150 0.849 0.803 0.948 0.867
CS [12] CVPR’24 0.030 0.110 0.100 0.080 0.905 0.906 0.903 0.905
ours 0.021 0.085 0.000 0.035 0.895 0.894 0.953 0.914
III-C Applications under Different Scenarios
III-C1 Multi-Agent Planning
Our pipeline can be useful in training models that learn to determine trajectories or actions for controlling one or multiple agents simultaneously. Furthermore, users can customize these scenarios by setting them under varying weather conditions, such as during a cloudy day or a rainy night. Another application is simulating the condition with emergency vehicles, enabling scenarios where ambulances, fire trucks, or police vehicles navigate without impediments, thereby enhancing the realism and utility of the training data. We provide examples in Fig. 3.
III-C2 Critical Scenario
A significant aspect of our scene generation capability focuses on critical scenarios, which are essential for testing the robustness of autonomous driving systems. Our framework can be prompted to create scenarios with potentially dangerous settings. For example, in Fig. 4(a), we can generate scenes where visibility is obstructed given input presented in the sub-caption, causing pedestrians and vehicles to be unseen by each other. Additionally, with the road conditions, we can create an unprotected left-turn scene with no traffic light and stop sign as presented in Fig. 4(b).
IV Experiments
IV-A Pipeline Setup
In the experimental setup, we utilize the CARLA simulator and the built-in maps from Town01 to Town07 to construct our road database. Roads lacking sufficient space for agent deployment are excluded. Our agent database accommodates nine distinct types of agents: ambulance, police car, firetruck, bus, truck, motorcycle, cyclist, car, and pedestrian. The simulation engine supports six actions: turning, lane changing, straight, stop, crossing road, and blocking, with denoting optional left or right directions. Additionally, as for position relative to the ego agent, our framework supports twelve directional positions, including eight basic compass directions and specific placements like left road, right road, opposite straight road, and near the ego destination. We illustrate the action and the relative position in Fig. 5. As for the LLM agent, we adopt the GPT-4o [23] to achieve analysis, road candidate retrieval, and agent planning.
IV-B Generation for Critical Scenarios
To assess the efficacy of our framework to train an autonomous agent, we examine our approach to critical scenarios with SafeBench [39], a benchmark for the safety evaluation of autonomous driving systems. Our pipeline creates scenes across three distinct challenging safety scenarios selected from ChatScene [12], aiming to conduct adversarial training to bolster the ego vehicle’s safe-driving skills. We prompt the LLM agent with the scenario descriptions and use the planning results from the LLM and our rendering interface to train the ego agents under a controlled policy using a soft-actor-critic model [40]. During training, the ego agent is controlled by a learnable policy model, and our engine controls other agents that create the scenario.
[…]
In this table, we compare our results with Learning-to-Collide (LC) [36], AdvSim (AS) [37], Adversarial Trajectory Optimization (AT) [38], and ChatScene (CS) [12]. Our findings indicate that our text-to-scene pipeline effectively trains agents to handle critical conditions, thereby enhancing safety and reducing collision risks. Notably, since our pipeline can provide diverse scenes during training, our agents are proficient at managing obstacles, lane changing, and yielding to oncoming cars. This can be observed by reducing the average collision rate from 0.08 to 0.035. However, reducing collisions may come at the cost of comfort (e.g., sudden swerving or hard braking), leading to slightly lower overall scores in certain scenarios. Overall, our approach effectively enables agents to learn from critical scenario experiences and enhances their safety capabilities.
Figure 5: Illustration of relative position (left) and the action (right).
TABLE II: Evaluation Results of Scene Diversity.
Metric Scenario Avg.
Normal Critical Conditional
Daily Traffic Intersection Pedestrian Crushing Blocking Agent Dangerous Cut-off Only having Two-wheel Vehicles Having Emergency Vehicles Rainy Weather
[…]
IV-C Ablation Study
The primary objective of our framework design is to provide a reliable pipeline to generate traffic scenes. Therefore, we evaluate the efficacy of different components by adopting ten distinct prompts under diverse scenarios. The scenarios encompass 3 normal, four critical, and three specific road conditions, such as the presence of traffic lights or the absence of crossroads or a stop sign on the road.
TABLE III: Analysis of pipeline design.
Quality Scenario Avg.
Normal Critical Conditional
[…]
w/o. analysis .917 .667 .833 .750 .750 .917 .833 .775
w./o analysis .917 1.00 .875 .750 1.00 .917 .925 .875
Scene Quality $\uparrow TM\uparrow$ $\uparrow TM\uparrow$
w/o. ranking 0.667 0.450 0.600 10.560
w./o ranking 0.867 0.750 0.800 0.800
IV-C1 Plan Quality
We assess plan quality by measuring the accuracy of planning results from the LLM agent using two metrics: Agent Accuracy (AA) and Road Accuracy (RA). AA evaluates the correctness of agent attributes, including position, action, and type. RA assesses the accuracy of the road conditions, such as signals, objects, and lane counts. As indicated in the first part of Table III, incorporating an analysis step into the reasoning process significantly enhances performance. This is because the analysis step provides the agent with more contextual information, leading to more precise decisions. Crucially, integrating an analysis step helps stabilize the number of agents, significantly mitigating the issue of generating unnecessary agents that substantially detract from the agent’s accuracy.
IV-C2 Scene Quality
To evaluate the impact of our road ranking strategy on scene quality, we use Text Matching (TM) to examine the correctness between the prompt and the generated scene. To ensure fairness and isolate the effect of the ranking strategy, identical road conditions and agent planning are maintained across evaluations, and each scenario is tested five times to calculate average outcomes. The results are detailed in the second part of Table III that the ranking approach consistently shows superior performance. The improvement is largely because, without ranking, the selection of roads is random, often disregarding crucial factors such as turn permissions or the adequacy of spawning points. Conversely, the road ranking strategy ensures the selection of roads that optimally align with the specified agent planning, thus enhancing the fidelity of the generated traffic scenes to the input prompts.
V Discussion
To further clarify our approach, we conduct additional experiments to discuss the diversity (Sec. V-A) and demonstrate an extra application in driving captioning (Sec. V-B).
V-A Diversity Test
A key claim of our approach is its ability to generate varied traffic scenes without reliance on fixed selections or predefined spawning points. We construct eight distinct scenarios under three main categories, normal, critical, and conditional, to validate the diversity of our generated results. To examine our pipeline’s ability to generate proper and diverse results, we provide no clues about the roads and agents within the text input. Each scenario is initiated with a text prompt formatted as: “Please create a scene for <scenario>” to generate traffic scenes.
We assess the diversity of these scenarios by generating each scene five times and compute the metrics, including Agent Diversity (AD) and Road Diversity (RD). These diversity metrics are calculated as the ratio of unique objects to the total number of objects. For agents, diversity considers variations in agent type, action, and relative position; for roads, diversity is assessed by unique road IDs (different directions on the same road are considered the same). Additionally, to verify the practicality and accuracy of our rendering interface, we include the metric Text Matching (TM), which evaluates whether the generated scenes accurately reflect the input prompts using a binary matched/unmatched criterion. The comprehensive results are detailed in Table II.
As the table indicates, our pipeline consistently provides a nearly unique road selection for each scenario. For the AD score, we observe lower scores in scenarios with specific types such as pedestrian crushing and emergency vehicles. This outcome is anticipated as more detailed agent descriptions naturally lead to similar planning results. Conversely, scenarios with less specific cues, such as daily traffic or intersection under normal scenarios, blocking agent under critical scenarios, or rainy weather under conditional scenarios, achieve higher diversity scores. These results demonstrate that our approach effectively leverages the LLM agent to generate diverse road and agent planning. Regarding the TM scores, despite lower scores in scenarios that involve precise timing (pedestrian crushing) or have restrictive conditions (only having two-wheel vehicles), our rendering interface accurately produce the majority of the scenarios, allowing us to obtain precise traffic scenes from text inputs.
V-B Driving Captioning
Since our pipeline provides images from the car’s front view, it can also be utilized for training driving captioning models for action reasoning under critical scenes. Captioning models, such as ADAPT [7] and BehaviorGPT4 [29], generate narrations of driving actions and their reasoning based on video input. However, as these models are trained on real-world datasets like BDD-X [41], they lack exposure to critical scenarios.
We illustrate an example in Fig. 6. In this case, the expected action is “turns left slowly,” with the reason of “giving way to a coming car.” However, the model fails to capture “slow” and only describes the left turn. Additionally, the generated reasoning is incorrect due to hallucination, further indicating that the original model does not effectively account for critical scenarios.
To improve the model’s critical reasoning capabilities, we sample 20 critical scenes for training and another 20 for evaluation from our pipeline. To ensure the collected scenes accurately reflect how the ego-agent responds to critical situations, we set its behavior to be cautious and resample whenever a collision occurs. Since the captions generated by our pipeline do not naturally follow a narration-with-reasoning structure, we use an LLM to reformat the scene descriptions to match the training format. For our experiments, we select the model from ADAPT [7], as it does not rely on closed-source models like GPT-4 and can be trained efficiently without excessive computational resources.
The results in Tab. IV are evaluated using standard captioning metrics, including BLEU [42], METEOR [6], and CIDEr [43], on our critical scenario data. Before fine-tuning, the model performs significantly worse on critical scenarios, particularly in reasoning. It struggles to recognize critical objects and fails to infer the appropriate actions. In contrast, after learning from a small set of examples, the model improves significantly in reasoning about actions, allowing it to generate more reliable descriptions for abnormal or critical scenes. Additionally, the descriptions of actions become more precise, leading to better captioning ability.
Figure 6: Illustration of predicted captions under critical scenario generation. N denotes narration, while R represents reasoning. The subject of narration is omitted as it always refers to the ego-car.
TABLE IV: Results of Driving Captioning on critical scenario.
Narration Reasoning
B M C B M C
original 4.8 13.5 15.2 0.0 10.0 18.4
finetuned 9.9 18.4 52.9 7.2 11.2 51.9
VI Failure Analysis
In this section, we analyze the failure cases observed in our experiments and discuss potential reasons and solutions.
VI-A Planning Failures
Since our framework is built upon a Large Language Model (LLM), the generated scenes highly depend on the LLM agent’s planning capabilities. Although we incorporate a reasoning stage to enhance the accuracy of both the agent and the road conditions, the LLM can still produce incorrect results in certain cases. One common issue is the LLM’s tendency to overlook alternative phrasing. For example, if the input prompt describes the scenario as “a car in front does not move for an infinite amount of time,” the LLM agent may fail to recognize this as a blocking action. Another challenge arises from hallucinations generated by the LLM agent. For instance, when instructed to generate two or more cars, the model frequently introduces a traffic light as a required signal for the road, even when it is not specified in the prompt. To mitigate these issues, one possible approach is to employ a reasoning LLM agent such as OpenAI-o1 [44] to provide better planning results. Additionally, fine-tuning an open-source LLM using expert demonstrations or preference alignment [45] can help reduce hallucinations and improve overall accuracy.
VI-B Retrieval Failures
Another source of failure arises from missing signals and objects within the CARLA environment. Since our framework relies on the built-in map, complex combinations of signals and objects and novel conditions cannot be retrieved. For instance, identifying a road that includes left, right, and straight directions while excluding traffic lights and stop signs is currently infeasible. This limitation exists because all all-directional intersections are designed to include either a traffic light or a stop sign. Additionally, while a speed limit sign displaying “70” is supported, requesting a speed limit sign of “200” is not. Similar limitations arise when users specify conditions that are not pre-defined in the database. To overcome these constraints, we plan to enhance support for generating non-existent signals, objects, and roads by directly creating OpenDRIVE code and integrating it into the map. This approach will allow users to customize various signals and objects, enabling more diverse scene generation.
VII Conclusion and Future Work
In this paper, we introduce a novel framework that leverages natural language descriptions for traffic scene generation facilitated by a Large Language Model. Our pipeline is able to understand scenario requirements effectively and find the best virtual environment locations for scene generation with diverse and customized options. Our pipeline starts by analyzing user input to decompose the requirements, performing road candidate retrieval, and planning detailed agents. To ensure the selected scene is suitable for an agent to act and be spawned, we adopt a road ranking strategy to match all the conditions. We provide experiments on diversity tests and adversarial training on the critical scenes to examine our pipeline ability to generate diverse and reliable scenes. Looking ahead, we plan to enhance the system’s practicality for agent and road generation to enable the creation of non-existent signals and objects and design a generated model that can transfer simulated scenes into a real-world video. This will further expand our framework’s ability to create a real and dynamic scene.